The functions f and g are defined by f(x) =1/(1-3x) and g(x) =log1/3(3x-2)-log3(x) respectively



1. Write down the sets Df (ehe domain of f) and Dg (the domain of g)



2. Solve the inequality f(x) > 2 for x\is in∈ Df


3. Solve the inequality f(x) ≥ 2 for x\is in∈ Dg


Hint: Use the change of base formula

1
Expert's answer
2022-01-04T11:27:06-0500

1.


"f(x)=\dfrac{1}{1-3x}"

"1-3x\not=0=>x\not=\dfrac{1}{3}"

"Df: (-\infin, \dfrac{1}{3})\cup(\dfrac{1}{3}, \infin)"

"g(x)=\log_{1/3}(3x-2)-\log_{3}(x)"

"\begin{cases}
 3x-2>0 \\
 x>0 
\end{cases}=>x>\dfrac{2}{3}"

"Dg: (\dfrac{2}{3}, \infin)"

2.


"f(x)>2, x\in Df"

"\dfrac{1}{1-3x}>2, x\in(-\infin, \dfrac{1}{3})\cup(\dfrac{1}{3}, \infin)"

"\dfrac{1}{1-3x}-2>0"

"\dfrac{1-2+6x}{1-3x}>0"

"\dfrac{6x-1}{3x-1}<0"

"\dfrac{1}{6}<x<\dfrac{1}{3}"

"x\in(\dfrac{1}{6},\dfrac{1}{3})"

3.


"g(x)\geq2, x\in Dg"

"\log_{1/3}(3x-2)-\log_{3}(x)\geq2, x\in(\dfrac{2}{3}, \infin)"

"\log_{1/3}(3x-2)=-\log_{3}(3x-2)"

"2=\log_{3}(9)"

"-\log_{3}(3x-2)-\log_{3}(x)\geq\log_{3}(9)"

"\log_{3}(9x(3x-2))\leq0"

"3>1=>y(x)=\log_{3}(x)\ increases"

Then


"9x(3x-2)\leq1, x\in(\dfrac{2}{3}, \infin)"

"27x^2-18x-1\leq0"

Let


"27x^2-18x-1=0"

"D=(-18)^2-4(27)(-1)=432"

"x=\dfrac{18\pm\sqrt{432}}{2(27)}=\dfrac{3\pm2\sqrt{3}}{9}=\dfrac{1}{3}\pm\dfrac{2\sqrt{3}}{9}"

"\begin{cases}
 \dfrac{1}{3}-\dfrac{2\sqrt{3}}{9}\leq x\leq \dfrac{1}{3}+\dfrac{2\sqrt{3}}{9} \\
 x>\dfrac{2}{3} 
\end{cases}"

"=>\dfrac{2}{3}<x\leq\dfrac{1}{3}+\dfrac{2\sqrt{3}}{9}"

"x\in\bigg(\dfrac{2}{3},\ \dfrac{1}{3}+\dfrac{2\sqrt{3}}{9}\bigg]"